OpenStudy (anonymous):

arithmetic sequence question! find t50 from: 3,9,15,21

OpenStudy (solomonzelman):

I got disconnected sorry

OpenStudy (solomonzelman):

So you are to find the 50th term of this sequence. You can see that you go ` + 6 ` each time to find the next term. You know that your 1st term is 3. For instance, if you were to find the 2nd term, you would add 6 once (and you would get 9) And if you were to find the 3rd term, you would add 6 twice (and you would get 12) So to find the 50th term, you would add number 6 49 times. 3 + 6 + 6 + 6 + \(\Large\color{black}{ ~^{49~~\rm{ times} } }\) •••••• instead of adding all 49, would it be easier to say, 3 + (6 × 49) ?

OpenStudy (solomonzelman):

Can you finish the calculations (for the 50th term) ?

OpenStudy (anonymous):

so t50 = 441?

OpenStudy (solomonzelman):

t50 = 3 + (6 × 49) (this is not 441)

OpenStudy (anonymous):

oh so i get it; you just plug in tn - a + d (n-1) ?

OpenStudy (anonymous):

oh whoops

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ t_n = t_1 + d(n-1) }\)

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ t_{50}=3 +6(50-1) }\)

OpenStudy (solomonzelman):

need more help, or good from here ?

OpenStudy (anonymous):

so then it's t50 = 9(49)? im confused...

OpenStudy (solomonzelman):

No, \(\large\color{midnightblue}{ t_{50}=3 +6(50-1) }\) \(\large\color{midnightblue}{ t_{50}=3 +6(49) }\) fisrt multiply 6×49 and only then add 3.

OpenStudy (anonymous):

oh okay so that's 297...thank you for your help! i appreciate it :D

OpenStudy (solomonzelman):

Anytime :)

OpenStudy (solomonzelman):

(I learned the sequences in trigonometry, not in Alg1..... )